OpenStudy (anonymous):

Please help? Fan + Medal Kennedy can paint 11 square feet in 2 minutes. Assuming she paints at a constant rate, choose the linear equation that represents the situation.

OpenStudy (anonymous):

A. y= (121/2)x B. y= 22x (I think this one) C. y= (11/2)x D. y= (2/11)x

OpenStudy (isaidavila):

Make a ratio 11/y=2/x y=11/2x

OpenStudy (anonymous):

Okay.

OpenStudy (isaidavila):

If y represent the feet painted and x is the time

OpenStudy (anonymous):

Thank you